# DTN:  Deep Multiple Task-specific Feature Interactions Network for Multi-Task Recommendation

### v3

@@ Abstract
Neural-based multi-task learning (MTL) has been successfully applied to many recommendation applications. However, these MTL models (e.g., MMoE, PLE) did not consider feature interaction during the optimization, which is crucial for capturing complex high-order features and has been widely used in ranking models for real-world recommender systems. Moreover, through feature importance analysis across various tasks in MTL, we have observed an interesting divergence phenomenon that the same feature can have significantly different importance across different tasks in MTL. To address these issues, we propose Deep Multiple Task-specific Feature Interactions Network (DTN) with a novel model structure design. DTN introduces multiple diversified task-specific feature interaction methods and task-sensitive network in MTL networks, enabling the model to learn task-specific diversified feature interaction representations, which improves the efficiency of joint representation learning in a general setup. We applied DTN to our company’s real-world E-commerce recommendation dataset, which consisted of over 6.3 billion samples, the results demonstrated that DTN significantly outperformed state-of-the-art MTL models. Moreover, during online evaluation of DTN in a large-scale E-commerce recommender system, we observed a 3.28% in clicks, a 3.10% increase in orders and a 2.70% increase in GMV (Gross Merchandise Value) compared to the state-of-the-art MTL models. Finally, extensive offline experiments conducted on public benchmark datasets demonstrate that DTN can be applied to various scenarios beyond recommendations, enhancing the performance of ranking models.
11footnotetext: Both authors contributed equally to this research. Corresponding author is Yaowen Bi. 22footnotetext: This author is the one who did the really hard work for Online Testing and Visualization Analysis.
Introduction
Recommender systems (RS) plays a crucial role in user applications by providing personalized recommendations for preferred candidates such as items, ads, videos, news, and more, based on user properties. In an industrial setting, RS needs to fulfill various objectives, such as Click-Through Rate (CTR), Add-To-Cart rate (ATC), Conversion Rate (CVR), and so on. Multi-task learning (MTL), which
Figure 1: (a) Serial Feature Interaction approach in MTL (SFM). (b) Task-specific Feature Interactions (TFI) MTL model.
@@ Introduction
involves training a single network to generate multiple predictions for different objectives, is well-suited to meet these requirements. Consequently, there is an increasing trend in applying MTL to recommender systems (RS) to simultaneously model multiple aspects, including CTR, CVR and other objectives. In fact, this approach has become the predominant method in major industry applications (Ma et al. 2018b; Caruana 1997; Ma et al. 2018a; Maziarz et al. 2019; Tang et al. 2020).
@@
MTL has been demonstrated to enhance learning efficiency by facilitating information sharing among the tasks (Caruana 1997). However, these MTL models (e.g., MMoE, PLE) did not consider feature interaction during the optimization. Through feature importance analysis across various tasks in MTL, we have observed that the same feature can have significantly different importance values across different tasks in MTL, which is called divergence phenomenon in this paper. Therefore, in MTL tasks it is essential to consider feature interactions to learn better representations.
@@
To evaluate the performance of DTN, we conduct extensive experiments on both real-world industrial recommender systems and major publicly available datasets, including census-income (mis 2000), AE (pengcheng Li et al. 2020) and Ali-CCP (Ma et al. 2018a). The experiment results demonstrate that DTN outperforms state-of-the-art MTL models across all datasets, and also significantly better than the serial approach of feature interaction in MTL. Besides, significant improvement of online metrics on a large-scale E-commerce recommender system in our company demonstrates the advantage of DTN in real-world recommendation applications.
@@ 2. $\bullet$
    A DTN model with a novel feature interaction approach within MTL and a task-sensitive network is proposed to enhance the efficiency of joint representation learning for each individual task. Beyond recommendation applications, DTN is versatile and can be applied to a variety of scenarios.
@@ 3. $\bullet$
Feature Interaction
Many deep learning based ranking models have been proposed in recent years and it is the key factor for most of these neural network based models to effectively model the feature interactions, such as Factorization Machines (FM) (Rendle 2010), DeepFM(Guo et al. 2017), xDeepFM(Lian et al. 2018), DCN(Wang et al. 2017), DCN-M(Wang et al. 2020), FiBiNET(Huang, Zhang, and Zhang 2019), MaskNet(Wang, She, and Zhang 2021) and Momonet (Zhang and Zhang 2023) etc.
Recently, MaskNet (Wang, She, and Zhang 2021) has been proposed to learn specific representation vectors for each feature considering its context. DHEN (Zhang et al. 2022) leverages the strengths of heterogeneous interaction modules to learn a hierarchy of interactions across different orders. MemoNet (Zhang and Zhang 2023) utilizes a multi-Hash Codebook Network (HCNet) as a memory mechanism to efficiently learn and memorize representations of cross features in CTR tasks. GDCN (Wang et al. 2023) employs a GCN with an information gate to dynamically filter next-order cross features, enhancing performance and stability in capturing high-order interactions.
@@
Multi-Task Learning in Recommender Systems
To better leverage various user behaviors, multi-task learning has been widely applied in recommender systems, leading to significant improvements, such as ESMM (Ma et al. 2018b), MMoE (Ma et al. 2018a) , Customized Gate Control (CGC) (Tang et al. 2020) and PLE (Tang et al. 2020).
The ESSM (Ma et al. 2018b) introduces CTR, CVR task and shares embedding parameters to enhance the accuracy of CVR task. MMoE (Ma et al. 2018a) integrates shared experts using different gating networks for each specific task. PLE (Tang et al. 2020) explicitly separates task-shared and task-specific experts and employs a progressive routing mechanism to gradually extract and segregate deeper semantic knowledge.
@@
Divergence Phenomenon of Feature Importance
@@
Permutation Feature Importance
Feature effectiveness plays a crucial role in complex systems, particularly in ranking models used in real-world industrial recommender systems. In these systems, the identification of the most important feature sets is vital as it enables us to select key features that can maximize online prediction performance with limited computational resources. Feature importance can be obtained through various methods, including Dropout (Chang, Rampasek, and Goldenberg 2017), FSCD (Ma et al. 2021), and Permutation Feature Importance (Breiman 2001). We utilize permutation feature importance to obtain feature importance, which is widely used in real-world ranking models.
@@
Table 1: Feature Importance of features for task CTR/CVR
Figure 2: The Divergence Phenomenon: same feature have significantly different feature importance across CTR, CVR task in MTL, such as $order\_cnt\_24h$ and $item\_ctr\_48h$.
Figure 3:   (A) Architecture overview of DTN, which utilizes Multiple Feature Interaction (MFI) module and Task-Sensitive Network (TSN) to learn task-specific feature interaction representations.  (B)  TSN takes into account the sequential nature of tasks and leverages the representations of preceding tasks more effectively.  (C)  Gating Network (GN) is employed to compute the weighted sum of feature interaction representations.
Divergence Phenomenon of Feature Importance in MTL
$FI^{k}(X_{i})=AUC^{k}_{base}-AUC^{k}_{p}(X_{i}),$ (2)
There are nearly 400 features in the ranking model of our company’s E-commerce Recommender System. We analyzed the importance of all these features for both click-through rate (CTR) and conversion rate (CVR) tasks in our multi-task learning (MTL) ranking model using Permutation Feature Importance. The results are shown in Table 1, where we used $FI^{k}(X_{i})$ and $Rank^{k}_{FI}(X_{i})$ to assess the importance of feature $X_{i}$ for task $k$. The ranking, denoted by $Rank^{k}_{FI}(X_{i})$, represents the descending order of the features based on their corresponding $FI^{k}(X_{i})$ values for each task k. We presented several examples to demonstrate the divergence phenomenon of feature importance in Multi-Task Learning (MTL) in Table 1.
We can observe that the feature $item\_price$ and $realtime\_discount$ (real-time discount ratio of the item) both exhibit high importance in both the CTR and CVR tasks. However, the feature $item\_ctr\_48h$ (ctr of the item in recent 48 hours) and $image\_ebd$ (pre-trained embedding of item cover) show high importance in CTR but low importance in CVR. Conversely, the feature $order\_cnt\_24h$ (user order count in the recent 24 hours) and $atc\_items\_50$ demonstrate high importance in CVR but low importance in CTR. Certainly, there are many other features that can demonstrate this divergence phenomenon of feature importance across different tasks in MTL.
@@
This phenomenon illustrates the divergence of feature importance, wherein the significance of the same feature can vary substantially across different tasks in MTL. This emphasizes the significance of acquiring task-specific feature interaction representations in MTL.
Modeling Approaches
@@
SFM: Serial Feature Interaction Approach in MTL
@@
TFI: Task-Specific Feature Interaction
@@
Models Interaction CTR ATC CVR
AUC RelaImp LogLoss AUC RelaImp LogLoss AUC RelaImp LogLoss
PLE - 0.7401 - 0.5513 0.7800 - 0.4834 0.8052 - 0.4516
GDCN 0.7412 +0.46% 0.5473 0.7801 +0.04% 0.4830 0.8067 +0.49% 0.4462
SFM MemoNet 0.7424 +0.96% 0.5429 0.7808 +0.29% 0.4805 0.8083 +1.02% 0.4404
MaskNet 0.7448 +1.96% 0.5341 0.7816 +0.57% 0.4777 0.8095 +1.41% 0.4361
GDCN 0.7424 +0.96% 0.5429 0.7814 +0.50% 0.4784 0.8078 +0.85% 0.4422
TFI MemoNet 0.7440 +1.62% 0.5371 0.7813 +0.46% 0.4788 0.8076 +0.79% 0.4429
MaskNet 0.7473 +3.00% 0.5249 0.7844 +1.57% 0.4678 0.8107 +1.80% 0.4318
DTN(w/o TSN) MFI 0.7481 +3.33% 0.5220 0.7845 +1.61% 0.4674 0.8143 +2.98% 0.4188
DTN MFI 0.7483 +3.42% 0.5219 0.7850 +1.79% 0.4637 0.8153 +3.31% 0.4160
DTN-trim MFI+trim 0.7482 +3.37% 0.5231 0.7847 +1.68% 0.4667 0.8147 +3.11% 0.4174
Table 2: Performance comparison on CTR, Add-To-Cart (ATC) and CVR Task in our company dateset. We currently employ three newest feature interaction methods in DTN, such as GDCN, MemoNet and MaskNet for illustration. DTN-trim is the trimmed version of DTN that selects high-importance feature interactions for each task based on gating weights in our company dataset, which will introduce in feature interaction selection section
DTN: Deep Multiple Task-specific Feature Interactions Network
@@
Multiple diversified Feature Interactions
@@
Task-Sensitive Network
@@
Moreover, in the context of E-commerce Recommender Systems, it is important to acknowledge the sequential nature of multiple tasks, such as the requirement of a user to click on an item before being able to place an order. Hence, for the CVR task, we multiply the output of the $MFI_{T}$ with the predicted score of the CTR task to get final representations from the CTR task, which allows us to consider the inter-task correlation and enhance the modeling of subsequent tasks, which we refer to as the Task-Sensitive Network (TSN).
Specifically, TSN for CVR task contains two parts of input: the predicted score of CTR task $pCTR$, and output vecotr of MFI modules $M=[M_{T}(x),M_{S}(x),M_{V}(x)]$. By multiplying the $pCTR$ with the relevant MFI module, we obtain $A_{O}=[pCTR\cdot M_{T}(x),M_{S}(x)]$, with each vector in $A_{O}$ denoted as $A_{O}^{i}$. For the CTR task, only the gating network (GN) is utilized since it doesn’t involve any pre-task, unlike the CVR task.
In order to effectively retain the task-specific information learned for each task, our approach utilizes two gates for each task: one GN for the task-specific MFI and another for the task-shared and other task’s MFI components. For CVR task, the task-shared and other task’s MFI components we calculate the output $S_{O}$ as follows:
$S_{O}=\sum_{i}^{m}g_{O}^{i}A_{O}^{i}$ (4)
@@
$S_{V}=\sum_{i}^{n}g_{V}^{i}M_{V}(x)$ (5)
where $n$ is the total number of FIM modules in $M_{V}(x)$, so we have the output $C_{V}$ for CVR task as: $C_{V}=Concat[S_{O},S_{V}]$.
And the tower network is a multi-layer network with width and depth as hyper-parameters. The prediction of CVR task is:
$y_{V}=t_{V}C_{V},$ (6)
where $t_{V}$ denotes the tower unit of task CVR.
Analogously, we have the prediction of CTR task is:
$y_{T}=t_{T}C_{T},$ (7)
where $t_{T}$ denotes the tower unit of task CTR. The only difference is that the CTR task doesn’t have a pre-task, so only the gating network (GN) is used for it.
@@
Experiment

### v4

@@ Abstract.
Neural-based multi-task learning (MTL) has been successfully applied to many recommendation applications. However, these MTL models (e.g., MMoE, PLE) did not consider feature interaction during the optimization, which is crucial for capturing complex high-order features and has been widely used in ranking models for real-world recommender systems. Moreover, through feature importance analysis across various tasks in MTL, we have observed an interesting divergence phenomenon that the same feature can have significantly different importance across different tasks in MTL. To address these issues, we propose Deep Multiple Task-specific Feature Interactions Network (DTN) with a novel model structure design. DTN introduces multiple diversified task-specific feature interaction methods and task-sensitive network in MTL networks, enabling the model to learn task-specific diversified feature interaction representations, which improves the efficiency of joint representation learning in a general setup. We applied DTN to Shopee’s real-world E-commerce recommendation dataset, which consisted of over 6.3 billion samples, the results demonstrated that DTN significantly outperformed state-of-the-art MTL models. Moreover, during online evaluation of DTN in a large-scale E-commerce recommender system, we observed a 3.28% in clicks, a 3.10% increase in orders and a 2.70% increase in GMV (Gross Merchandise Value) compared to the state-of-the-art MTL models. Finally, extensive offline experiments conducted on public benchmark datasets demonstrate that DTN can be applied to various scenarios beyond recommendations, enhancing the performance of ranking models.
Recommender System, Multi-task Learning, Feature Interaction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August, 2026; TBAisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrieval
1. Introduction
@@
Recommender systems (RS) plays a crucial role in user applications by providing personalized recommendations for preferred candidates such as items, ads, videos, news, and more, based on user properties. In an industrial setting, RS needs to fulfill various objectives, such as Click-Through Rate (CTR), Add-To-Cart rate (ATC), Conversion Rate (CVR), and so on. Multi-task learning (MTL), which involves training a single network to generate multiple predictions for different objectives, is well-suited to meet these requirements. Consequently, there is an increasing trend in applying MTL to recommender systems (RS) to simultaneously model multiple aspects, including CTR, CVR and other objectives. In fact, this approach has become the predominant method in major industry applications (Ma et al., 2018a; Caruana, 1997; Ma et al., 2018b; Maziarz et al., 2019; Tang et al., 2020).
@@
MTL has been demonstrated to enhance learning efficiency by facilitating information sharing among the tasks (Caruana, 1997). However, these MTL models (e.g., MMoE, PLE) did not consider feature interaction during the optimization. Through feature importance analysis across various tasks in MTL, we have observed that the same feature can have significantly different importance values across different tasks in MTL, which is called divergence phenomenon in this paper. Therefore, in MTL tasks it is essential to consider feature interactions to learn better representations.
@@
To evaluate the performance of DTN, we conduct extensive experiments on both real-world industrial recommender systems and major publicly available datasets, including census-income (mis, 2000), AE (pengcheng Li et al., 2020) and Ali-CCP (Ma et al., 2018b). The experiment results demonstrate that DTN outperforms state-of-the-art MTL models across all datasets, and also significantly better than the serial approach of feature interaction in MTL. Besides, significant improvement of online metrics on a large-scale E-commerce recommender system in Shopee demonstrates the advantage of DTN in real-world recommendation applications.
@@ 2. $\bullet$
    A DTN model with a novel feature interaction approach within MTL and a task-sensitive network is proposed to enhance the efficiency of joint representation learning for each individual task. Meanwhile, the MFI module can also be used independently for single-task learning, such as CTR prediction task. Beyond recommendation applications, DTN is versatile and can be applied to a variety of scenarios.
@@ 2. Related Work
Figure 1. (a) Serial Feature Interaction approach in MTL (SFM). (b) Task-specific Feature Interactions (TFI) MTL model.
2.1. Feature Interaction
Many deep learning based ranking models have been proposed in recent years and it is the key factor for most of these neural network based models to effectively model the feature interactions, such as Factorization Machines (FM) (Rendle, 2010), DeepFM(Guo et al., 2017), xDeepFM(Lian et al., 2018), DCN(Wang et al., 2017), DCN-M(Wang et al., 2020), FiBiNET(Huang et al., 2019), MaskNet(Wang et al., 2021) and Momonet (Zhang and Zhang, 2023) etc.
Recently, MaskNet (Wang et al., 2021) has been proposed to learn specific representation vectors for each feature considering its context. DHEN (Zhang et al., 2022) leverages the strengths of heterogeneous interaction modules to learn a hierarchy of interactions across different orders. MemoNet (Zhang and Zhang, 2023) utilizes a multi-Hash Codebook Network (HCNet) as a memory mechanism to efficiently learn and memorize representations of cross features in CTR tasks. GDCN (Wang et al., 2023) employs a GCN with an information gate to dynamically filter next-order cross features, enhancing performance and stability in capturing high-order interactions.
@@ 2.1. Feature Interaction
2.2. Multi-Task Learning in Recommender Systems
To better leverage various user behaviors, multi-task learning has been widely applied in recommender systems, leading to significant improvements, such as ESMM (Ma et al., 2018a), MMoE (Ma et al., 2018b) , Customized Gate Control (CGC) (Tang et al., 2020) and PLE (Tang et al., 2020).
The ESSM (Ma et al., 2018a) introduces CTR, CVR task and shares embedding parameters to enhance the accuracy of CVR task. MMoE (Ma et al., 2018b) integrates shared experts using different gating networks for each specific task. PLE (Tang et al., 2020) explicitly separates task-shared and task-specific experts and employs a progressive routing mechanism to gradually extract and segregate deeper semantic knowledge.
@@ 2.2. Multi-Task Learning in Recommender Systems
3. Divergence Phenomenon of Feature Importance
@@
3.1. Permutation Feature Importance
Feature effectiveness plays a crucial role in complex systems, particularly in ranking models used in real-world industrial recommender systems. In these systems, the identification of the most important feature sets is vital as it enables us to select key features that can maximize online prediction performance with limited computational resources. Feature importance can be obtained through various methods, including Dropout (Chang et al., 2017), FSCD (Ma et al., 2021), and Permutation Feature Importance (Breiman, 2001). We utilize permutation feature importance to obtain feature importance, which is widely used in real-world ranking models.
@@ 3.1. Permutation Feature Importance
3.2. Divergence Phenomenon of Feature Importance in MTL
(2) $FI^{k}(X_{i})=AUC^{k}_{base}-AUC^{k}_{p}(X_{i}),$
There are nearly 400 features in the ranking model of Shopee’s E-commerce Recommender System. We analyzed the importance of all these features for both click-through rate (CTR) and conversion rate (CVR) tasks in our multi-task learning (MTL) ranking model using Permutation Feature Importance. The results are shown in Table 1, where we used $FI^{k}(X_{i})$ and $Rank^{k}_{FI}(X_{i})$ to assess the importance of feature $X_{i}$ for task $k$. The ranking, denoted by $Rank^{k}_{FI}(X_{i})$, represents the descending order of the features based on their corresponding $FI^{k}(X_{i})$ values for each task k. We presented several examples to demonstrate the divergence phenomenon of feature importance in Multi-Task Learning (MTL) in Table 1.
Table 1. Feature Importance (FI) of features for task CTR/CVR
@@ 3.2. Divergence Phenomenon of Feature Importance in MTL
We can observe that the features $item\_price$ and $realtime\_discount$ (real-time discount ratio of each item) exhibit high importance in both the CTR and CVR tasks. However, the feature $item\_ctr\_48h$ (ctr of the item in recent 48 hours) and $image\_ebd$ (pre-trained embedding of item cover) show high importance in CTR but low importance in CVR. Conversely, the feature $order\_cnt\_24h$ (user order count in the recent 24 hours) and $atc\_items\_50$ demonstrate high importance in CVR but low importance in CTR. Certainly, there are many other features that can demonstrate this divergence phenomenon of feature importance across different tasks in MTL.
@@
This phenomenon illustrates the divergence of feature importance, wherein the significance of the same feature can vary substantially across different tasks in MTL. This divergence extends to feature interactions, which are constructed by combining basic features to generate high-order representations. Their effectiveness and importance also vary across tasks in MTL. Therefore, introducing task-specific feature interaction mechanisms is essential, as it enables the learning of task-specific interaction representations, thereby improving model performance and generalization. In Section 5.4, we also offer a comprehensive analysis of the variations in result representations among different feature interaction methods across multiple tasks in MTL.
Figure 2. The Divergence Phenomenon: same feature have significantly different feature importance across CTR, CVR task in MTL, such as $order\_cnt\_24h$ and $item\_ctr\_48h$.
4. Modeling Approaches
@@
4.1. SFM: Serial Feature Interaction Approach in MTL
@@
4.2. TFI: Task-Specific Feature Interaction
@@
Figure 3.   (A) Architecture overview of DTN, which utilizes Multiple Feature Interaction (MFI) module and Task-Sensitive Network (TSN) to learn task-specific feature interaction representations.  (B)  TSN takes into account the sequential nature of tasks and leverages the representations of preceding tasks more effectively.  (C)  Gating Network (GN) is employed to compute the weighted sum of feature interaction representations.
4.3. DTN: Deep Multiple Task-specific Feature Interactions Network
@@
4.3.1. Multiple diversified Feature Interactions
@@
4.3.2. Task-Sensitive Network
@@
Moreover, it is important to acknowledge the sequential nature of multiple tasks in MTL, such as the requirement for users to click on an item before being able to place an order, as well as the need for users to click on a video in video recommendation scenarios in order to have the duration of play.
Therefore, for the subsequent task $V$, we introduce a more explicit signal by multiplying the prediction score of the preceding task $T$ with the output of $MFI_{T}$ to obtain the final representation of the preceding task $T$. This allows us to consider the interdependencies between tasks and enhance the modeling of subsequent tasks. We refer to this approach as Task-Sensitive Network (TSN).
Specifically, TSN for task $V$ contains two parts of input: the predicted score of $T$ task $pred\_T$, and output vecotr of MFI modules $M=[M_{T}(x),M_{S}(x),M_{V}(x)]$. By multiplying the $pred\_T$ with the relevant MFI module, we obtain $A_{O}=[pred\_T\cdot M_{T}(x),M_{S}(x)]$, with each vector in $A_{O}$ denoted as $A_{O}^{i}$. For the $T$ task, only the gating network (GN) is utilized since it doesn’t involve any pre-task, unlike the $V$ task.
In order to effectively retain the task-specific information learned for each task, our approach utilizes two gates for each task: one GN for the task-specific MFI and another for the task-shared and other task’s MFI components. For task $V$, the task-shared and other task’s MFI components we calculate the output $S_{O}$ as follows:
(4) $S_{O}=\sum_{i}^{m}g_{O}^{i}A_{O}^{i}$
@@
(5) $S_{V}=\sum_{i}^{n}g_{V}^{i}M_{V}(x)$
where $n$ is the total number of FIM modules in $M_{V}(x)$, so we have the output $C_{V}$ for task $V$ as: $C_{V}=Concat[S_{O},S_{V}]$.
And the tower network is a multi-layer network with width and depth as hyper-parameters. The prediction of task $V$ is:
(6) $y_{V}=t_{V}C_{V},$
where $t_{V}$ denotes the tower unit of task $V$.
Analogously, we have the prediction of $T$ task is:
(7) $y_{T}=t_{T}C_{T},$
where $t_{T}$ denotes the tower unit of task $T$. The only difference is that the task $T$ doesn’t have a pre-task, so only the gating network (GN) is used for it.
@@
Table 2. Performance comparison on CTR, Add-To-Cart (ATC) and CVR Task in Shopee dateset. We currently employ three newest feature interaction methods in DTN, such as GDCN, MemoNet and MaskNet for illustration. DTN-trim is the trimmed version of DTN that selects high-importance feature interactions for each task based on gating weights in Shopee dataset, which will introduce in section 5.3 .
Models Feature Interactions CTR ATC CVR
AUC RelaImpr LogLoss AUC RelaImpr LogLoss AUC RelaImpr LogLoss
Shared-Bottom - 0.7341 - 0.5631 0.7759 - 0.4931 0.8029 - 0.4627
MMoE - 0.7363 +0.93% 0.5571 0.7786 +0.97% 0.4881 0.8041 +0.39% 0.4572
PLE - 0.7401 +2.56% 0.5513 0.7800 +1.48% 0.4834 0.8052 +0.75% 0.4516
GDCN 0.7412 +3.03% 0.5473 0.7801 +1.52% 0.4830 0.8067 +1.25% 0.4462
SFM MemoNet 0.7424 +3.54% 0.5429 0.7808 +1.77% 0.4805 0.8083 +1.78% 0.4404
MaskNet 0.7448 +4.57% 0.5341 0.7816 +2.06% 0.4777 0.8095 +2.17% 0.4361
GDCN 0.7424 +3.54% 0.5429 0.7814 +1.99% 0.4784 0.8078 +1.61% 0.4422
TFI MemoNet 0.7440 +4.22% 0.5371 0.7813 +1.95% 0.4788 0.8076 +1.55% 0.4429
MaskNet 0.7473 +5.63% 0.5249 0.7844 +3.08% 0.4678 0.8107 +1.38% 0.4318
DTN(w/o TSN) MFI 0.7481 +5.98% 0.5220 0.7845 +3.11% 0.4674 0.8143 +3.92% 0.4188
DTN MFI 0.7483 +6.06% 0.5219 0.7850 +3.29% 0.4637 0.8153 +4.09% 0.4160
DTN-trim MFI+trim 0.7482 +6.02% 0.5231 0.7847 +3.19% 0.4667 0.8147 +3.89% 0.4174
5. Experiment